55. The length if a rectangle is 1 more than 3 times the width. If the width is increased by 2 and the length is doubled, the new perimeter is 3 less than 5 times the original length. Find the original dimensions.
Answer by dabanfield(803)

The length if a rectangle is 1 more than 3 times the width. If the width is increased by 2 and the length is doubled, the new perimeter is 3 less than 5 times the original length. Find the original dimensions.
Let l be the length and w the width of the rectangle:
1.) l = 3*w + 1
perimeter = 2*l + 2*w so:

2.) 2*(l*2) + 2*(w+2) = 5*l - 3
Substitute 3*w + 1 for l in equation 2.) and solve for w. Then calculate l = 3*w + 1